OpenStudy (anonymous):

Solve -8b+k=-3 for k. How would I solve this?

jimthompson5910 (jim_thompson5910):

\[\Large -8b+k=-3 \] \[\Large k=-3+8b \]

jimthompson5910 (jim_thompson5910):

We're basically subtracting 8b on the left side, so we "undo" this by adding 8b to both sides

jimthompson5910 (jim_thompson5910):

this will isolate k

OpenStudy (anonymous):

How come you don't add 8b

OpenStudy (anonymous):

he did add 8b

jimthompson5910 (jim_thompson5910):

on the second line I'm adding 8b to both sides, on the left side it's not shown since -8b+8b = 0b = 0 which goes away (which is what we want)